Find the probability that of the next 200 children born, (a) less than 40% will be boys, (b) between 43% and 57% will be girls, (c) more than 54% will be boys. Assume equal probabilities for births of boys and girls

1
Expert's answer
2021-11-29T15:57:36-0500

If "X" is a binomial random variable with mean "\u03bc = np" and variance "\sigma^2=np(1-p)," then the limiting form of the distribution of

"Z=\dfrac{X-np}{\sqrt{np(1-p)}},"

as "n\to \infin," is the standard normal distribution "n(z; 0, 1)."

In practice, the approximation is adequate provided that both "np\geq10" and "np(1-p)\geq 10."

a) "n=200, p=0.5, \mu=np=200(0.5)=100,"

"\sigma^2=np(1-p)=200(0.5)(1-0.5)=50"


"P(X<80)\approx P(X<80-0.5)"

"\approx P(Z<-2.8991378)"

"\approx0.001871"

b) "n=200, p=0.5, \mu=np=200(0.5)=100,"

"\sigma^2=np(1-p)=200(0.5)(1-0.5)=50"


"P(86<X<114)=P(X<114)-P(X\leq 86)"

"\approx P(X<114-0.5)-P(X<86+0.5)"

"=P(Z<\dfrac{113.5-100}{\sqrt{50}})-P(Z<\dfrac{86.5-100}{\sqrt{50}})"

"\approx P(Z<1.90918831)-P(Z<-1.90918831)"

"\approx0.9718811-0.0281189\approx0.943762"

c) "n=200, p=0.5, \mu=np=200(0.5)=100,"

"\sigma^2=np(1-p)=200(0.5)(1-0.5)=50"


"P(X>108)=1-P(X\leq 108)"

"\approx 1-P(X<108+0.5)"

"=1-P(Z<\dfrac{108.5-100}{\sqrt{50}})"

"\approx1-P(Z<1.20208153)"

"\approx 0.114666"